The ages of three sisters are three even consecutive integers. If the sum of the 1st, four times the 2nd, and twice the 3rd is 86, what are the three ages?
1
Expert's answer
2013-01-21T11:09:49-0500
Let age of the first sister x, then for second x+2, for the third x+4 we have x+4(x+2)+2(x+4)=86

x+4x+8+2x+8=86
7x+16=86
7x=70
x=10
so the ages are 10,12 and 14